1.         50 billion barrels of oil were consumed in the world in 2018. There is a projected 2,400 billion barrels of oil remaining under ground.

a.    How much oil would be consumed in year 2048 (30 years later) under the following annual growth rates:

            1.         5%

            2.         3%

            3.         no growth

            4.         - 2%

b.   How much TOTAL oil would be consumed over 30 years under the same growth rate scenarios [5%, 3%, 0%, -2%]
Answer #1

50*(105/100)*(103/100)*(98/100)=52.99 approx 53 billion barral oil will be used in 2048

Total oil consumed=50*(105/100)+50*(105/100)*(103/100)+50*(105/100)*(103/100)+50*(105/100)*(103/100)*(98/100)= 52.5+54.75+54.75+52.99= 214.99 or approximately 215 bilion barrels of oil will be used